# A bottleneck model with shared autonomous vehicles: Scale economies and price regulations

### v1

@@ Abstract
This study examines how scale economies in the operation of shared autonomous vehicles (SAVs) affect the efficiency of a transportation system where SAVs coexist with normal vehicles (NVs). We develop a bottleneck model where commuters choose their departure times and mode of travel between SAVs and NVs, and analyze equilibria under three SAV-fare scenarios: marginal-cost pricing, average-cost pricing, and unregulated monopoly pricing. Marginal-cost pricing reduces commuting costs but results in financial deficits for the service provider. Average-cost pricing ensures financial sustainability but has contrasting effects depending on the timing of implementation due to the existence of multiple equilibria: when implemented too early, it discourages adoption of SAVs and increases commuting costs; when introduced after SAV adoption reaches the monopoly equilibrium level, it promotes high adoption and achieves substantial cost reductions without a deficit. We also show that expanding road capacity may increase commuting costs under average-cost pricing, demonstrating the Downs–Thomson paradox in transportation systems with SAVs. We next examine two optimal policies that improve social cost, including the operator’s profit: the first-best policy that combines marginal-cost pricing with congestion tolls, and the second-best policy that relies on fare regulation alone. Our analysis shows that these policies can limit excessive adoption by discouraging overuse of SAVs. This suggests that promoting SAV adoption does not always lower social cost.
@@ 1 Introduction
One significant challenge arising from the presence of scale economies is the formation of natural monopolies. When fixed costs are high and scale economies are strong, a single provider can supply the services demanded by users at a lower cost than any combination of two or more providers. This cost advantage enables the provider to dominate the market, naturally leading to a monopoly. Natural monopolies are a common feature of public transport systems and shared mobility services (Hörcher and Tirachini, 2021). Without regulation, a monopolistic provider prioritizes profit maximization, resulting in higher fares for commuters and inefficiencies in service provision. Therefore, effective fare regulations, such as marginal cost pricing or average cost pricing, are essential to mitigate the inefficiencies (Varian and Melitz, 2024).
@@
Another important challenge is the emergence of a feedback loop between service adoption and operational efficiency. As more users adopt a shared service, the average cost falls, which improves service quality and attracts even more users. Conversely, a decline in adoption increases average costs, degrading service quality and further discouraging usage. Such feedback mechanisms can lead to multiple equilibria, where the system may settle into a high- or low-adoption equilibrium, depending on initial conditions.
@@
These feedback mechanisms can also give rise to paradoxical system responses, where seemingly beneficial interventions lead to unintended consequences. A notable example is the classic Downs–Thomson paradox in public transport (Downs, 1962; Thomson, 1977). In that context, improvements in road infrastructure reduce public transport ridership, raise its average cost, and degrade its quality. A similar dynamic may emerge even as SAVs and NVs share the same road space: as road capacity expands, some commuters may shift to NVs, reducing SAV utilization and increasing average costs. This, in turn, leads to a deterioration in service quality and higher total commuting costs.
@@
Recent theoretical work has examined the impacts of SAVs on transportation systems. Among these, the bottleneck model has become a central analytical tool because it effectively captures key features of peak-period traffic congestion, which are essential for evaluating commuter behavior under limited road capacity. Using this framework, many studies have focused on two main effects of SAVs: the capacity effect, which increases road throughput by reducing headways, and the value-of-time (VOT) effect, which reflects the time savings from allowing commuters to engage in other activities during in-vehicle travel. These effects have been extensively analyzed in the literature on shared and autonomous mobility systems (e.g., van den Berg and Verhoef, 2016; Tian et al., 2019; Yu et al., 2022b). [^1] Other modeling approaches have also been employed to analyze VOT and capacity effects during peak-period congestion. Notably, Dantsuji and Takayama (2024) adopt a bathtub model, which captures capacity drops that are not incorporated in bottleneck models, to evaluate how AVs influence congestion dynamics through these effects.
@@
Although these effects have been extensively studied, relatively little attention has been paid to scale economies, which are a defining feature of shared mobility services. This gap is important because scale economies directly influence how fare structures affect user adoption, operational efficiency, and overall system performance. Therefore, important interactions among pricing, adoption behavior, and system efficiency remain insufficiently understood in the current literature.
@@
This study shows the influence of scale economies on the efficiency of transport systems with SAVs. To this end, we develop a bottleneck model that incorporates SAVs to examine how scale economies and natural monopoly characteristics affect commuters’ mode choices and traffic congestion. Using this model, we analyze three fare-setting scenarios: marginal cost pricing, average cost pricing, and unregulated monopoly pricing. In addition, we evaluate two transport policies designed to achieve an efficient transportation system. The first-best policy combines marginal cost pricing with vehicle-specific, time-varying congestion tolls to address inefficiencies from both natural monopoly and congestion. The second-best policy assumes that congestion tolls are not feasible and focuses on fare regulation alone.
Focusing first on fare-setting scenarios, we examine the effects of marginal cost pricing and average cost pricing. Marginal cost pricing reduces commuting costs by encouraging SAV adoption. However, because fares are set equal to marginal costs, it fails to cover fixed costs and leads to financial deficits for the provider. Average cost pricing ensures financial sustainability by covering both fixed and variable costs. Our analysis shows that its effectiveness depends critically on the timing of implementation. If introduced during the early stage of SAV deployment, it discourages adoption and leads to higher commuting costs compared to the case without regulation. In contrast, allowing monopolistic operation in the initial phase and applying average cost pricing at a later stage can promote adoption, improve service efficiency, and reduce commuting costs. This outcome reflects a feedback loop: limited adoption spreads fixed costs over fewer trips, raising average costs and further discouraging usage.
@@
We also evaluate how road capacity expansion interacts with these fare-setting scenarios. We show that, under average cost pricing, capacity expansion may unintentionally increase total commuting costs. This happens because improved road conditions encourage commuters to switch back to NVs, which lowers SAV utilization and raises the average cost of providing the service. This outcome demonstrates the Downs–Thomson paradox in transport systems with SAVs.
We further demonstrate that the optimal policy implications depend critically on the relative strength of the two key characteristics of SAVs: the VOT effect and the capacity effect. We first consider the first-best policy, which aims to minimize social cost. When the capacity effect is stronger than the VOT effect, this policy promotes SAV adoption and results in all commuters being better off compared to any equilibrium without policy. By contrast, when the VOT effect dominates, the first-best policy may reduce SAV adoption, since eliminating queuing delays diminishes the relative advantage of SAVs.
We next consider the second-best policy, where congestion tolls cannot be implemented and social cost is minimized through fare regulation alone. Our analysis reveals that, compared to the equilibrium without policy, the second-best policy can reduce SAV usage when the VOT effect dominates the capacity effect. This outcome arises because higher SAV adoption exacerbates queuing delays through the VOT effect, while the capacity effect is insufficient to alleviate the resulting congestion. To prevent excessive usage, the fare under the second-best policy may be set above the average cost to intentionally discourage SAV adoption.
@@
Furthermore, the results of the second-best policy analysis imply that marginal cost pricing does not always minimize social cost, even in the absence of policy intervention. When the VOT effect dominates the capacity effect, both average cost pricing and monopoly pricing can result in lower social costs than under marginal cost pricing by limiting excessive SAV usage. In particular, monopoly pricing may yield the lowest social cost, as profit-driven fare setting can inadvertently discourage excessive adoption of SAVs. These findings highlight the importance of aligning fare policy with SAV technology characteristics and caution against universally applying marginal cost pricing in the absence of congestion tolls.
@@
Related Literature: [^2] For a comprehensive review of the economics of public transport systems, including fare regulation and market structure, see Hörcher and Tirachini (2021). For a broad survey of the expected impacts, modeling approaches, and policy considerations surrounding SAVs, see Narayanan et al. (2020). The implications of scale economies in public transport systems have long been studied in the economics and transport literature. Many theoretical studies have examined mode choice between private vehicles and public transport systems characterized by scale economies (e.g., Tabuchi, 1993; Danielis and Marcucci, 2002). A number of theoretical studies have examined the Downs–Thomson paradox in multimodal transport systems that explicitly account for scale economies in rail transit services, including the effects of fare structures and operational objectives (e.g., Arnott and Yan, 2000; Basso and Jara-Díaz, 2012; Bell and Wichiensin, 2012; Zhang et al., 2014; Wang et al., 2019). In the context of bus transit systems, several theoretical studies have also incorporated scale economies (e.g., Cantarella et al., 2015; Li and Yang, 2016; Li et al., 2018; Pandey and Lehe, 2024).
@@
A growing body of research has explored the implications of autonomous vehicles (AVs) for urban transportation systems, focusing on their potential to change commuter behavior and alleviate congestion. Building on Vickrey (1969)’s bottleneck model, numerous studies have examined AV impacts on departure time choice and mode competition. Early works considered the VOT effect and the capacity effect due to AVs (e.g., van den Berg and Verhoef, 2016), while later studies incorporated SAVs, parking constraints, and policy instruments such as dedicated lanes and reservation systems (e.g., Tian et al., 2019; Lamotte et al., 2017; Li et al., 2022). Other extensions introduced in-vehicle activity utility (Pudāne, 2020; Yu et al., 2022a; Wu and Li, 2023) and the competition and cooperation between different car manufacturers that may provide both NVs and AVs (Yu et al., 2022b). These studies have significantly advanced our understanding of AV-induced behavioral shifts, but they do not explicitly address provider-side issues such as scale economies or fare regulation, which are central to SAV services.333 Recent studies have considered scale economies in shared mobility systems, particularly on-demand ridepooling services (Fielbaum et al., 2023; Fielbaum and Pudāne, 2024). These studies focus on how demand affects system-level cost structures through matching efficiency, waiting times, and detour lengths. However, they do not use bottleneck models or examine the policy implications of fare regulation or road capacity expansion. Our study complements this literature by extending a bottleneck model to examine how scale economies interact with fare regulations and pricing policies in shaping commuter behavior and system efficiency.
This study contributes to the literature by explicitly modeling the supply-side structure of SAV services, focusing on the implications of scale economies and the resulting natural monopoly. We examine how different fare-setting scenarios, such as marginal cost pricing, average cost pricing, and monopoly pricing, affect user adoption, congestion, and overall system efficiency. Our framework also reveals that expanding road capacity can have unintended consequences under average cost pricing, such as reducing SAV adoption and increasing commuting costs. We investigate first-best and second-best policy outcomes and show that these policies can limit excessive adoption by discouraging overuse of SAVs when the VOT effect dominates. This implies that promoting SAV usage is not universally welfare-improving.
@@ 2.1 Network, commuters, and transport modes
A fixed number of commuters travel to the CBD along this road. They are treated as a continuum with total mass $N$. Commuters use either NVs or SAVs. The number of NV commuters is denoted by $N_{n}$, and that of SAV commuters is denoted by $N_{a}$. NV commuters own their vehicles, incurring fixed costs such as parking, denoted by $F_{n}$. A service provider operates SAVs and requires SAV commuters to pay a fare $p$.
Following van den Berg and Verhoef (2016), we assume that SAVs have two effects: a capacity effect and a VOT effect. Bottleneck capacity increases to $\mu/\kappa$ $(0<\kappa<1)$ when SAVs pass through since they can safely drive closer together than NVs. For VOT, SAV commuters experience a reduction since they can engage in other activities during in-vehicle time. Let $\alpha$ denote the VOT for NV commuters. The VOT for SAV commuters is $\theta\cdot\alpha$, where $\theta$ $(0<\theta<1)$ is the VOT reduction parameter. Without loss of generality, we set $\alpha=1$.
@@ 2.2 Behavior of commuters and the service provider
Each commuter chooses their departure time and transport mode to minimize the commuting cost, which consists of free-flow travel time, queuing delay, schedule delay, and mode-specific costs. Considering the capacity effects of SAVs, we define $q(t)$ by the queuing delay for commuters arriving at the destination (CBD) at time $t$.
@@
$c_{a}(t)=\theta\left\{t_{f}+q(t)\right\}+s(t)+p,$ (3)
@@
which is strictly decreasing in $N_{a}$; hence the service provider benefits from scale economies. Scale economies often lead to a natural monopoly, allowing the provider to set fares that maximize profit. This may result in an inefficient transportation system in which commuters face high commuting costs. To mitigate such distortions, a public authority commonly introduces price regulation.
In this study, we compare the following three fare-setting scenarios to evaluate the impact of price regulation on the NV-SAV transportation system.
@@
    Marginal-cost (MC) pricing. The fare is set equal to the marginal operating cost:
@@
    Natural monopoly. The service provider sets the fare to maximize profit $\pi$ given the demand function $N_{a}(p)$. Here, $N_{a}(p)$ denotes the demand for SAV service at fare $p$; its analytical form will be derived later as the equilibrium outcome of commuters’ departure-time and mode choices. The fare $p$ satisfies the first-order condition of profit maximization:

### v2

@@ Abstract
This study examines how scale economies in the operation of shared autonomous vehicles (SAVs) affect the efficiency of a transportation system where SAVs coexist with normal vehicles (NVs). We develop a bottleneck model where commuters choose their departure times and mode of travel between SAVs and NVs, and analyze equilibria under three SAV fare-setting scenarios: marginal cost pricing, average cost pricing, and unregulated monopoly pricing. Marginal cost pricing reduces commuting costs but results in financial deficits for the service provider. Average cost pricing ensures financial sustainability but has contrasting effects depending on the timing of implementation due to the existence of multiple equilibria: when implemented too early, it discourages adoption of SAVs and increases commuting costs; when introduced after SAV adoption reaches the monopoly equilibrium level, it promotes high adoption and achieves substantial cost reductions without a deficit. We also show that expanding road capacity may increase commuting costs under average cost pricing, demonstrating the Downs–Thomson paradox in transportation systems with SAVs. We next examine two optimal policies that improve social cost, including the operator’s profit: the first-best policy that combines marginal cost pricing with congestion tolls, and the second-best policy that relies on fare regulation alone. Our analysis shows that these policies can limit excessive adoption by discouraging overuse of SAVs. This suggests that promoting SAV adoption does not always reduce social cost.
@@ 1.1 Background
Among these challenges, one important implication of scale economies is the potential emergence of natural monopoly. When fixed costs are high and scale economies are strong, a single provider can supply the services demanded by users at a lower cost than any combination of two or more providers. This cost advantage enables the provider to dominate the market, naturally leading to a monopoly. Natural monopolies are a common feature of public transport systems and shared mobility services (Hörcher and Tirachini, 2021). Without regulation, a monopolistic provider prioritizes profit maximization, resulting in higher fares for commuters and inefficiencies in service provision. Therefore, effective fare regulations, such as marginal cost pricing or average cost pricing, are essential to mitigate the inefficiencies (Varian and Melitz, 2024).
@@
In addition to these implications, scale economies also introduce a feedback loop between service adoption and operational efficiency. As more users adopt a shared service, the average cost falls, which improves service quality and attracts even more users. Conversely, a decline in adoption increases average costs, degrading service quality and further discouraging usage. Such feedback mechanisms can lead to multiple equilibria, where the system may settle into a high- or low-adoption equilibrium, depending on initial conditions.
@@
The mechanisms can, in turn, give rise to paradoxical system responses, where seemingly beneficial interventions lead to unintended consequences. A notable example is the classic Downs–Thomson paradox in public transport (Downs, 1962; Thomson, 1977). In that context, improvements in road infrastructure reduce public transport ridership, raise its average cost, and degrade its quality. A similar dynamic may emerge even as SAVs and NVs share the same road space: as road capacity expands, some commuters may shift to NVs, reducing SAV utilization and increasing average costs. This, in turn, leads to a deterioration in service quality and higher total commuting costs.
@@
Recent theoretical work has examined the impacts of SAVs on transportation systems. Among these, the bottleneck model has become a central analytical tool because it effectively captures key features of peak-period traffic congestion, which are essential for evaluating commuter behavior under limited road capacity. Using this framework, many studies have focused on two main effects of SAVs: the capacity effect, which increases road throughput by reducing headways, and the value-of-time (VOT) effect, which reflects the time savings from allowing commuters to engage in other activities during in-vehicle travel. These effects have been extensively analyzed in the literature on shared and autonomous mobility systems (e.g., van den Berg and Verhoef, 2016; Tian et al., 2019b; Yu et al., 2022b). [^1] Other modeling approaches have also been employed to analyze VOT and capacity effects during peak-period congestion. Notably, Dantsuji and Takayama (2024) adopt a bathtub model, which captures capacity drops that are not incorporated in bottleneck models, to evaluate how AVs influence congestion dynamics through these effects.
@@
Despite this growing body of work, relatively little attention has been paid to scale economies, which are a defining feature of shared mobility services. This gap is important because scale economies directly influence how fare structures affect user adoption, operational efficiency, and overall system performance. Therefore, important interactions among pricing, adoption behavior, and system efficiency remain insufficiently understood in the current literature.
1.2 Contribution and overview
@@
This study shows the influence of scale economies on the efficiency of transport systems with SAVs. To this end, we develop a bottleneck model that incorporates SAVs to examine how scale economies and natural monopoly characteristics affect commuters’ mode choices and traffic congestion. Using this model, we analyze three SAV fare-setting scenarios: marginal cost pricing, average cost pricing, and unregulated monopoly pricing. In addition, we evaluate two transport policies designed to achieve an efficient transportation system. The first-best policy combines marginal cost pricing with vehicle-specific, time-varying congestion tolls to address inefficiencies from both natural monopoly and congestion. The second-best policy assumes that congestion tolls are not feasible and focuses on fare regulation alone.
The analysis reveals that fare-setting rules shape SAV adoption through distinct feedback mechanisms induced by scale economies. Marginal cost pricing lowers commuting costs by encouraging SAV usage, but it cannot sustain service provision in the presence of fixed costs, resulting in financial deficits. Average cost pricing ensures cost recovery, but its impacts on system performance depend critically on the level of SAV usage at the time of implementation. If average cost pricing is implemented when SAV adoption remains limited, the resulting high fares suppress usage and increase commuting costs. By contrast, when fares are initially unregulated and the provider initially sets profit-maximizing fares, subsequent implementation of average cost pricing can sustain high adoption levels, improve service efficiency, and reduce commuting costs. This contrast arises from a feedback loop induced by scale economies, whereby adoption levels and average costs mutually reinforce each other.
@@
In addition, road capacity expansion can interact with fare regulation in counterintuitive ways. Under average cost pricing, expanding road capacity may unintentionally increase total commuting costs. This occurs because improved road conditions encourage commuters to switch back to NVs, which reduces SAV utilization and raises the average cost of providing the service. As a result, the commuting cost faced by users increases, demonstrating the Downs–Thomson paradox in transport systems with SAVs.
@@
The policy implications depend critically on the relative strength of two key characteristics of SAVs: the capacity effect and the VOT effect. Under the first-best policy, SAV adoption is promoted when the capacity effect dominates, whereas it may be reduced when the VOT effect is stronger. Under the second-best policy, where policy intervention is limited to fare regulation, welfare-improving outcomes may require limiting SAV usage through higher fares when the VOT effect dominates (i.e., excessive SAV adoption can arise without fare regulation). A key implication of the policy analysis is that marginal cost pricing does not universally minimize social cost in the absence of congestion tolls. When the VOT effect dominates the capacity effect, average cost pricing or even unregulated monopoly pricing can yield lower social costs by discouraging excessive SAV adoption. Overall, these results underscore the importance of aligning fare policy with the underlying technological characteristics of SAVs.
1.3 Related Literature
@@
The implications of scale economies in public transport systems have long been studied in the economics and transport literature. [^2] For a comprehensive review of the economics of public transport systems, including fare regulation and market structure, see Hörcher and Tirachini (2021). For a broad survey of the expected impacts, modeling approaches, and policy considerations surrounding SAVs, see Narayanan et al. (2020). Many theoretical studies have examined mode choice between private vehicles and public transport systems characterized by scale economies (e.g., Tabuchi, 1993; Danielis and Marcucci, 2002). A number of theoretical studies have examined the Downs–Thomson paradox in multimodal transport systems with rail transit (e.g., Arnott and Yan, 2000; Basso and Jara-Díaz, 2012; Bell and Wichiensin, 2012; Zhang et al., 2014; Wang et al., 2019). In the context of bus transit systems, several theoretical studies have also incorporated scale economies (e.g., Cantarella et al., 2015; Li and Yang, 2016; Li et al., 2018; Pandey and Lehe, 2024).
@@
A growing body of research has explored the implications of autonomous vehicles (AVs) for urban transportation systems, focusing on their potential to change commuter behavior and alleviate congestion. Building on Vickrey (1969)’s bottleneck model, numerous studies have examined AV impacts on departure time choice and mode competition. Early works considered the VOT effect and the capacity effect due to AVs (e.g., van den Berg and Verhoef, 2016), while later studies incorporated SAVs, parking constraints, and policy instruments such as dedicated lanes and reservation systems (e.g., Tian et al., 2019b; Lamotte et al., 2017; Li et al., 2022). Other extensions introduced in-vehicle activity utility (Pudāne, 2020; Yu et al., 2022a; Wu and Li, 2023) and the competition and cooperation between different car manufacturers that may provide both NVs and AVs (Yu et al., 2022b). These studies have significantly advanced our understanding of AV-induced behavioral shifts, but they do not explicitly address provider-side issues such as scale economies or fare regulation, which are central to SAV services. [^3] Recent studies have considered scale economies in shared mobility systems, particularly on-demand ridepooling services (Fielbaum et al., 2023; Fielbaum and Pudāne, 2024). These studies focus on how demand affects system-level cost structures through matching efficiency, waiting times, and detour lengths. However, they do not use bottleneck models or examine the policy implications of fare regulation or road capacity expansion. Our study complements this literature by extending a bottleneck model to examine how scale economies interact with fare regulations and pricing policies in shaping commuter behavior and system efficiency.
This study develops a bottleneck model of shared autonomous vehicles that explicitly incorporates scale economies and natural monopoly. By embedding a fixed cost into a multimodal departure time choice framework, the model enables a unified analysis of marginal cost pricing, average cost pricing, and unregulated monopoly pricing for SAV services. This framework allows us to examine how fare-setting rules shape SAV adoption, congestion patterns, and system performance through feedback between usage levels and cost recovery. By providing a transparent setting in which pricing regimes and infrastructure changes can be directly compared within a unified framework, the study offers new insights into the design of fare regulation for shared autonomous mobility under congestion.
@@ 2.1 Network, commuters, and transport modes
A fixed number of commuters travel to the CBD along this road. They are treated as a continuum with total mass $N$. Commuters use either NVs or SAVs. The number of NV commuters is denoted by $N_{n}$, and that of SAV commuters is denoted by $N_{a}$. NV commuters own their vehicles and incur a fixed cost term $F_{n}$. This term includes costs related to parking, such as parking fees. By contrast, SAVs are operated by a service provider. Thus, SAV commuters do not incur such costs, and instead pay a fare $p$ to the provider. In addition, using an SAV involves a pickup waiting time. For tractability, we represent its monetary equivalent by a constant $w$, in line with previous studies (e.g., Ma and Zhang, 2017; Tian et al., 2019a). [^4] In reality, pickup waiting time may depend on the provider’s fleet operations (e.g., fleet sizing, dispatching, and repositioning). This study abstracts from these supply-side operational decisions and treats the pickup waiting time as an exogenous constant to keep the focus on the interaction between scale economies and commuters’ demand-side choices. Extensions that endogenize fleet operations are briefly discussed in Section 6 as directions for future work.
Following van den Berg and Verhoef (2016), we assume that SAVs have two effects: a capacity effect and a VOT effect. For the capacity effect, we assume that the bottleneck capacity becomes $\mu/\kappa$ when an SAV passes through it. Here, $\kappa$ is an aggregate parameter that captures the overall capacity impact of SAV-related technologies, including connectivity and automation. [^5] In mixed traffic, the capacity effects may depend on the penetration rate of connected-vehicle technology. However, the literature shows that heterogeneity in VOT can induce equilibrium sorting that generates temporally concentrated SAV flows in a bottleneck setting. We therefore assume that connectivity operates primarily within these flows, and summarize the resulting capacity effect using the reduced-form parameter $\kappa$. In this study, we focus on the case $0<\kappa<1$, which represents a forward-looking scenario in which mature automation and connectivity allow SAVs to operate safely at shorter headways, thereby making capacity improvements feasible. For the VOT effect, SAV commuters experience a reduction because they can engage in other activities during in-vehicle time. Let $\alpha$ denote the VOT for NV commuters for in-vehicle time. The VOT for SAV commuters is $\theta\cdot\alpha$, where $\theta$ $(0<\theta<1)$ is the VOT reduction parameter. Without loss of generality, we set $\alpha=1$.
@@ 2.2 Behavior of commuters and the service provider
Each commuter chooses their departure time and transport mode to minimize the commuting cost, which consists of free-flow travel time, queuing delay, schedule delay, and mode-specific costs. We define $q(t)$ as the bottleneck queuing delay for commuters arriving at the destination (CBD) at time $t$.
@@
$c_{a}(t)=\theta\left\{t_{f}+q(t)\right\}+s(t)+p+w,$ (3)
@@
which is strictly decreasing in $N_{a}$; hence the service provider benefits from scale economies. Scale economies often lead to a natural monopoly, under which the provider may set the profit-maximizing fares. This may result in an inefficient transportation system in which commuters face high commuting costs. To mitigate such distortions, a public authority commonly introduces fare regulation.
In this study, we compare the following three fare-setting scenarios to evaluate the impact of fare regulation on the NV–SAV transportation system.
@@
    Marginal cost (MC) pricing. The fare is set equal to the marginal operating cost:
@@
    Unregulated monopoly pricing. The service provider sets the fare to maximize profit $\pi$ given the demand function $N_{a}(p)$. Here, $N_{a}(p)$ denotes the demand for SAV service at fare $p$; its analytical form will be derived later as the equilibrium outcome of commuters’ departure-time and mode choices. The fare $p$ satisfies the first-order condition of profit maximization: